OpenStudy (anonymous):

Consider a toy car that moves back and forth on a long, straight track for 4 minutes in such a way that the function s(t)=96t-84t^2+28t^3-3t^4 describes how far the car is to the right of the starting point, in centimeters, after t minutes. When is the toy car accelerating the fastest to the right? When is the toy car accelerating fastest to the left?

OpenStudy (anonymous):

I am having a hard time with trying to optimize this problem :(

OpenStudy (anonymous):

translate this as "find the maximum value of the second derivative of \([0,4]\)

OpenStudy (anonymous):

*of and i guess also the minimum of the second derivative

OpenStudy (anonymous):

i get \[-12 (3 t^2-14 t+14)\]

OpenStudy (anonymous):

Ok so what I have so far is the first, second, and third derivative of the equation and I calculated t=2.33 for the third derivative. I then put that number into the second derivative as well as 0 and 4 and I had 2.33 being the most positive and 0 being the most negative otu of the three values.

OpenStudy (anonymous):

max is at the vertex min is at 0 i think

OpenStudy (anonymous):

so do i need the third derivative at all?

OpenStudy (anonymous):

yeah you got 0 as the most negative, so that looks good the vertex is \(\frac{14}{6}=\frac{7}{3}\) you can use it if you like, but you don't need calculus to find the max and min or a quadratic check the endpoints and the vertex

OpenStudy (anonymous):

of course you will get the vertex by taking the derivative, setting it equal to zero and solve \[f(t)=at^2+bt+c\] \[f'(t)=2at+b\] \[2at+b=0\iff t=-\frac{b}{2a}\] same thing

OpenStudy (anonymous):

ok let me just change it quickly...

OpenStudy (anonymous):

no need, i think you got the same answer, just used a decimal instead of my fraction

OpenStudy (anonymous):

oh i erased my work already :( this is a little easier for me though

OpenStudy (anonymous):

since \(\frac{7}{3}=2\frac{1}{3}=2.333...\)

OpenStudy (anonymous):

the first derivative is so long :( i have so far... t=8+7t^2-t^3/4

OpenStudy (anonymous):

Ok I got it back so I have t=2.33 having the most positive s''(t) value and t=4 having the most negative value

OpenStudy (anonymous):

sorry i mean t=0 as most negative

OpenStudy (anonymous):

so that would mean...the car accelerates the fastest to the right at t=2.33 and the car accelerates fastest to the left at t=0?

OpenStudy (anonymous):

or at least i think it is right lol